OpenStudy (anonymous):

Which line shows the first error in the solution? A number divided by 3 is four more than 11. What is the number? line A: x/3 + 4 = -11 line B: x/3 + 4 - 4 = -11 - 4 line C: x/3 = -15 line D: x/3 times 3 = -15 times 3 line E: x= -45

OpenStudy (michele_laino):

I think the first line: so line A.

OpenStudy (michele_laino):

because, line A should be like this: x/3+4=11

OpenStudy (anonymous):

Thank you!

OpenStudy (michele_laino):

sorry: x/3=11+4

OpenStudy (ageta):

one is wrong because the sentence says 4 less then x/3 The equation they give is x/3+4=-11 When it is supposed to be x/3-4=-11

OpenStudy (anonymous):

\[x \div3=11+4 \]

sammixboo (sammixboo):

A number (n) divided by 3 (\(n\div 3\)) is four more than 11 (\(n\div 3~=~4~+11\)) than eleven

OpenStudy (ageta):

When it is supposed to be x/3-4=-11

OpenStudy (anonymous):

It should be that one I think or is it this one: \[x \div3=11-3\]

OpenStudy (anonymous):

It is line A.

OpenStudy (ageta):

yes the answer would be A

OpenStudy (anonymous):

Right? I mean . . . is it + or -?

sammixboo (sammixboo):

is basically means equal to, and since it says 4 more than 11, it would be +

sammixboo (sammixboo):



OpenStudy (anonymous):

\[x \div3=11+4\]